OpenStudy (anonymous):

Simplify the expression. 8x5+2x7 -------- 2(7-4)

OpenStudy (cwrw238):

you need to find the GCF of 8x^5 and 2x^7 this is the highest term which divides into these 2 terms can u do this?

OpenStudy (cwrw238):

first find GCF of 2 and 8 - that is 2 now how about x^5 and x^7?

OpenStudy (cwrw238):

its x^5 because x^7 / x^5 = x^2 and x^5 / x^5 = 1 so GCF is 2x^5 now you can factor the top of the fraction by taking 2x^5 out r u ok with that?

OpenStudy (anonymous):

nope